# Using cognitive models to reveal value trade-offs in language models

### v3

@@ Abstract
Value trade-offs are an integral part of human decision-making and language use, however, current tools for interpreting such dynamic and multi-faceted notions of values in LLMs are limited. In cognitive science, so-called “cognitive models” provide formal accounts of such trade-offs in humans, by modeling the weighting of a speaker’s competing utility functions in choosing an action or utterance. Here we use a leading cognitive model of polite speech to systematically evaluate value trade-offs in two encompassing model settings: degrees of reasoning “effort” in frontier black-box models, and RL post-training dynamics of open-source models. Our results highlight patterns of higher informational utility than social utility in reasoning models’ default behavior, and demonstrate that these patterns shift in predictable ways when models are prompted to prioritize certain goals over others. Our findings from LLMs’ training dynamics suggest large shifts in utility values early on in training with persistent effects of the choice of base model and pretraining data, compared to feedback dataset or alignment method. Our framework offers a flexible tool for probing value trade-offs across diverse model types, providing insights for generating hypotheses about other social behaviors such as sycophancy and for shaping training regimes that better control trade-offs between values during model development.
@@ 1 Introduction
A large body of work in cognitive science has formalized pragmatic communication in humans as a family of recursive probabilistic generative models, known as Rational Speech Acts (RSA) models (Frank & Goodman, 2012; Goodman & Frank, 2016).
@@
This class of cognitive models includes a pragmatic speaker that chooses what to say by balancing a mixture of goals (including being informative, but also various other affective, relational, and persuasive goals), and a pragmatic listener that interprets the speaker’s utterances and actions by taking into account such possible goals (e.g. Kao et al., 2014; Kaufmann et al., 2024; Barnett et al., 2022; Carcassi & Franke, 2023; Sumers et al., 2024).
@@ 1.1 Contributions
We apply this tool to a variety of closed and open-source large language models (see Appendix Table 1), and demonstrate the relevance of a structured probabilistic model of cognitive processes as a distinctive method for model interpretation. Our closed-source model suite consists of three families of frontier models across three values of reasoning budget (none, low, and medium), with analyses of the effects of prompt-based manipulations that simulate different “goals” a speaker can have (to be informative, social, or both). Our open-source model suite is designed to disentangle the roles of model family, feedback dataset, and alignment method in the RL post-training process. We infer the parameters of the cognitive model over training checkpoints for a total of 8 unique configurations of these aspects.
Our results highlight patterns of higher informational utility than social utility in reasoning models’ default behavior, and demonstrate that these patterns shift in predictable ways when models are prompted to prioritize certain goals over others. Further, models’ training dynamics over the alignment process reveal that the largest shifts in utility values happen within the first quarter of training. Still, it appears that the choice of base model and pretraining data may have an outsized impact on the resulting weighting of utilities compared to the choice of feedback dataset or alignment method. Taken together, our findings suggest that this method is responsive to diverse aspects of the rapidly evolving LLM landscape: our tool provides opportunities for forming fine-grained hypotheses about other high-level behavioral concepts, understanding the extent of training needed to achieve particular values, and shaping recipes for higher-order reasoning and alignment capabilities.
@@ 2 Cognitive model
The utterance utility $U_{\text{total}}$ consists of three components that trade off according to a mixture parameter $\boldsymbol{\omega}$ of the pragmatic speaker $S_{2}$. The informational utility $U_{\textrm{inf}}(u;s)$ is formalized as $\log P_{L_{1}}(s|u)$, namely the degree to which a pragmatic listener $L_{1}$ infers the true state intended by the speaker. The social utility $U_{\textrm{soc}}(u)$ is formalized as $\mathbb{E}_{P_{L_{1}}(s|u)}[V(s)]$, capturing the extent to which a specific utterance by expectation induces social values for the listener $L_{1}$. The presentational utility $U_{\textrm{pre}}(u;\phi)$ is grounded on the pragmatic listener $L_{1}$’s inference about a first-order pragmatic speaker $S_{1}$, who solely trades off information goal and social goal. Mathematically, the presentational utility can be formalized as $\log P_{L_{1}}(\phi|u)$. This quantity captures the extent to which a pragmatic listener $L_{1}$ infers a specific value trade-off $\phi$ under their internal model of a first-order pragmatic speaker $S_{1}$, where $P_{L_{1}}(s,\phi|u)\propto P_{S_{1}}(u|s,\phi)P(s)P(\phi)$. In other words, $\phi$ is a trade-off that the speaker $S_{2}$ wants to project towards a lower-order pragmatic listener $L_{1}$. The utterance distributions of the first-order pragmatic speaker $S_{1}$ is as follows:
@@
The informativeness and the expected social value of an utterance $u$ are both a function of how the literal listener $L_{0}$ interprets utterances $P_{L_{0}}(s|u)$, which is grounded out on the literal semantics [[u]](s)[\![u]\!](s) with a prior over the states $s$ likely to be communicated, i.e. $P_{L_{0}}(s|u)\propto[\![u]\!](s)\cdot P(s)$. For simplicity, the mapping from true state $s$ (i.e. the speaker’s actual assessment of the listener’s creation, specified in terms of the number of stars they would give it; see Section 4.1) to its perceived social value, $V(s)$, is assumed to be an identity function.
@@ 3 Language model evaluation suites
We design two model suites for evaluation that cover a range of characteristics that are thought to have implications for LLMs’ ability to capture human-like value trade-offs (see Appendix Table 1). In the closed-source setting, we aim to understand the behavioral tendencies of widely-used black-box models and how their reasoning-optimized variants might be adapting LLM behaviors in everyday contexts where value alignment is critical (cf. Zhou et al., 2025; Huang et al., 2025; Jiang et al., 2025). In the open-source setting, we seek to understand which factors influence model behavior after preference fine-tuning by systematically evaluating the effects of base model family, preference dataset, and alignment algorithm on the resulting value trade-offs.
Closed-source model suite We evaluate three degrees of reasoning in Anthropic, Google, and OpenAI’s models: a) models that do not explicitly use any additional chain-of-thought reasoning (Claude-Sonnet-3.7, Gemini-Flash-2.0, and ChatGPT-4o), and b) the low and medium effort reasoning modes of their reasoning counterparts (Claude-Sonnet-3.7, Gemini-2.5-Flash, o4-mini). For the Gemini and o4 models, these effort levels can be specified directly by the parameters “low” and “medium”, but for Claude-Sonnet-3.7, which instead uses a specific token count, we map these values to 1k tokens and 8k tokens, respectively, following the values indicated in the Gemini API documentation.
Open-source model suite We consider all unique configurations of two 7B parameter base models (Qwen2.5-Instruct (Yang et al., 2024) and Llama-3.1-Instruct (Grattafiori et al., 2024), two feedback datasets (UltraFeedback and Anthropic HH-RLHF  (Bai et al., 2022a), and two learning algorithms (DPO (Rafailov et al., 2024) and PPO). These elements have been shown to have characteristics that shape downstream behavior, however the influence of each of these factors on learned value trade-offs remains unclear. For each configuration (8 total), we initialize from an instruction-tuned model, perform one epoch of supervised fine-tuning (SFT) on the ‘chosen’ responses, and follow with one epoch of preference optimization using either DPO or PPO (implemented using OpenRLHF (Hu et al., 2024)) with ArmoRM (Wang et al., 2024) as the reward model. We evaluate each model’s behavior across evenly spaced checkpoints throughout the preference fine-tuning stage to trace the evolution of alignment and value trade-offs (see Section C.1 for full hyperparameter details).
@@ 4.1 Experimental vignettes
We provide models with the same set of vignettes given to human participants in Yoon et al. (2020), which describe socially sensitive situations in which a speaker must convey their judgement of a listener’s creation (e.g. a poem, presentation, cake, etc.). The speaker’s actual opinion, or true state $s$, is expressed on a scale from 1 to 5 stars, where 1 is the lowest or most negative opinion, and 5 is the highest. [^1] We deviate from the original paper’s 0-3 heart scale to provide LLMs with a scale that is most natural to their training data, particularly online reviews. We find that this 1-5 star scale captures the semantic range of the available utterance options better than the original 0-3 scale. We present models with the set of eight utterance options $u$ (four descriptor words and their negations) in a multiple choice format:
@@ 4.2 Inferring cognitive model parameters
We implemented the inference model in Stan (Carpenter et al., 2017), a probabilistic programming language, and used the default Hamiltonian Monte Carlo implemented in Stan (No-U-Turn sampler, Hoffman et al. (2014)) to perform approximate inference of model parameters (see Section C.2 for further implementation details).
@@ 5.2 Closed-source model suite
Figure 3 shows the results of fitting the responses of Anthropic, Gemini, and OpenAI’s language models across three reasoning budgets (none, low, and medium) to the second-order speaker model. We establish that our inferred parameter values generalize to a held-out test split of the closed-source model data that includes all combinations of reasoning budget, vignette framing, communicative goal, and model family via a posterior predictive check, and find that the average MSE from our inferred parameters is significantly lower than that of randomly sampled parameter values from a reasonable prior ($\mu_{\textrm{inferred}}=0.03$, $\mu_{\textrm{random}}=0.06$; z=12.49z=-12.49; $p<0.001$; see Appendix Figure 6).
We begin with analyzing models’ default behavior in the absence of any explicit communicative goals (black lines). The parameter values of $\phi$ for the first-order speaker in $S_{2}$ measures the relative mixture of informativeness and social utility that a speaker $S_{2}$ wishes the other person to infer about them. We find that across model families, reasoning variants display higher $\phi$ values—a higher projected informational utility than social utility—than their non-reasoning counterparts. A linear mixed-effects model predicting the posterior mean $\phi$ from degrees of reasoning effect222model formula: phi $\sim$ reasoning_effort + (1|llm_family) + (1|framing) (reference level: no reasoning) with random intercepts of model family and vignette framing suggested a significant effect of both low and medium reasoning effort compared to the no-reasoning counterpart for default model behaviors ($\beta_{\text{low}}=0.228,t=6.338,p<.001$; $\beta_{\text{medium}}=0.211,t=5.846,p<.001$). The difference of the inferred $\phi$ among models of low and medium reasoning effort was not significant ($p=0.627$).
@@
5.2.1 Manipulating communicative goals
@@ 5.2.2 Signatures of sycophancy
Figure 5: Open-source LLM results. Inferred values of informational, social, and presentational utilities $\boldsymbol{\omega}$ (purple), and projected mixture of informational and social utilities $\phi$ (magenta), according to a cognitive model for LLMs’ training checkpoints across the RLHF process. Line variants indicate different combinations of base model and feedback dataset; rows = alignment method. Error bars indicate 95% high density region averaged across results from three framing manipulations. We find the largest shifts in values within the first quarter of training, with persistent effects of the choice of base model and pretraining data, compared to feedback dataset or alignment method.
@@ 5.3 Open-source model suite
Figure 5 shows the training dynamics of two base open-source LLMs, Qwen2.5-7B-Instruct (lighter) and Llama-3.1-8B-Instruct (darker), aligned to the UltraFeedback (dashed line) and Anthropic HH-RLHF (solid line) datasets, via DPO (top row) and PPO (bottom row). Across the different inferred parameters, we observe a number of consistent patterns within combinations of model and dataset. Across both PPO and DPO and the two feedback datasets, Qwen-instruct shows a higher $\omega_{\text{inf}}$, but lower weighting of $\omega_{\text{pre}}$ than Llama-instruct. The differences between the models’ weighting of social utility $\omega_{\text{soc}}$ are less pronounced, but still present, with Qwen-instruct generally converging to a lower weighting of social utility than Llama-instruct. The projected weighting towards informational utility in Qwen-instruct’s $\phi$, as well as its higher $\omega_{\text{inf}}$ compared to Llama-instruct aligns with prior work highlighting Qwen’s superior performance in mathematical and reasoning tasks compared to Llama (Gandhi et al., 2025; Zeng et al., 2025).
@@
For most of the inferred parameters, we do not observe significant qualitative differences in the training dynamic patterns resulting from PPO vs. DPO alignment methods. However, for the parameter $\phi$, PPO does appear to pull all four model and feedback dataset configurations to a similar mean value (approx. 0.7). In contrast, in the DPO case, Qwen-instruct appears to quickly converge to a greater weighting of informational utility, with $\phi$ almost equal to 1 in the case of alignment to both feedback datasets, which Llama-instruct shows more of a balance towards social utility (though it is still primarily information-leaning).
In general, we see that the largest shifts in utility values across all four parameters happen within the first quarter of training, consistent with earlier findings on rapid adaptation during RL post-training in mathematical domains (Zhao et al., 2025). While such prior work has emphasized the significance of the base model and its pretraining data Itzhak et al. (2025), our use of a shared supervised fine-tuning (SFT) stage on the same preference datasets across all models may attenuate these differences. Moreover, the relatively minor distinction between PPO and DPO in our results may be partly due to training both methods for only a single epoch, and the fact that the Armo-RM reward model used in PPO, was trained on subsets of the same UltraFeedback and Anthropic HH-RLHF datasets, further reducing divergence between the two approaches.

### v4

# arXiv:2506.20666v4[cs.CL] 02 Mar 2026

@@ Abstract
Value trade-offs are an integral part of human decision-making and language use, however, current tools for interpreting such dynamic and multi-faceted notions of values in language models are limited. In cognitive science, so-called “cognitive models” provide formal accounts of such trade-offs in humans, by modeling the weighting of a speaker’s competing utility functions in choosing an action or utterance. Here, we show that a leading cognitive model of polite speech can be used to systematically evaluate alignment-relevant trade-offs in language models via two encompassing settings: degrees of reasoning “effort” and system prompt manipulations in closed-source frontier models, and RL post-training dynamics of open-source models. Our results show that LLMs’ behavioral profiles under the cognitive model a) shift predictably when they are prompted to prioritize certain goals, b) are amplified by a small reasoning budget, and c) can be used to diagnose other social behaviors such as sycophancy. Our findings from LLMs’ post-training dynamics reveal large shifts in values early on in training and persistent effects of the choice of base model and pretraining data, compared to feedback dataset or alignment method. Our framework offers a flexible tool for probing behavioral profiles across diverse model types and gaining insights for shaping training regimes that better control trade-offs between values during model development.
@@ 1 Introduction
A large body of work in cognitive science has formalized pragmatic communication in humans as a family of recursive probabilistic generative models, known as Rational Speech Acts (RSA) models (Frank & Goodman, 2012; Goodman & Frank, 2016).This class of cognitive models includes a pragmatic speaker that chooses what to say by balancing a mixture of goals (including being informative, but also various other affective, relational, and persuasive goals), and a pragmatic listener that interprets the speaker’s utterances and actions by taking into account such possible goals (e.g. Kao et al., 2014; Kaufmann et al., 2024; Barnett et al., 2022; Carcassi & Franke, 2023; Sumers et al., 2024).
@@ 1.1 Contributions
We apply this tool to a variety of closed and open-source large language models (see Appendix Table 1), and demonstrate the relevance of a structured probabilistic model of cognitive processes as a distinctive method for model interpretation. Our closed-source model suite consists of three families of frontier models across three values of reasoning budget (none, low, and medium), with analyses of the effects of prompt-based manipulations that simulate different “goals” a speaker can have (to be informative, social, or both). Our open-source model suite is designed to disentangle the roles of model family, feedback dataset, and alignment method in the RL post-training process. We infer the parameters of the cognitive model over training checkpoints for a total of 8 unique configurations of these aspects [^1] Code and data for our analyses are available at https://github.com/skmur/many-wolves.
Our results show that LLMs’ behavioral profiles under the cognitive model a) shift predictably when they are prompted to prioritize certain goals, b) are amplified by a small reasoning budget, and c) can be used to diagnose other social behaviors such as sycophancy. Further, models’ training dynamics over the alignment process reveal that the largest shifts in utility values happen within the first quarter of training. Still, it appears that the choice of base model and pretraining data may have an outsized impact on the resulting weighting of utilities compared to the choice of feedback dataset or alignment method. Taken together, our findings suggest that this method is responsive to diverse aspects of the rapidly evolving LLM landscape: our tool provides opportunities for forming fine-grained hypotheses about other high-level behavioral concepts, understanding the extent of training needed to achieve particular values, and shaping recipes for higher-order reasoning and alignment capabilities.
@@ 2 Cognitive model
The utterance utility $U_{\text{total}}$ consists of three components that trade off according to a mixture parameter $\boldsymbol{\omega}$ of the pragmatic speaker $S_{2}$. The informational utility $U_{\textrm{inf}}(u;s)$ is formalized as $\log P_{L_{1}}(s|u)$, namely the degree to which a pragmatic listener $L_{1}$ infers the true state intended by the speaker. The social utility $U_{\textrm{soc}}(u)$ is formalized as $\mathbb{E}_{P_{L_{1}}(s|u)}[V(s)]$, capturing the extent to which a specific utterance by expectation induces social values for the listener $L_{1}$. For simplicity, the mapping from true state $s$ (i.e. the speaker’s actual assessment of the listener’s creation, specified in terms of the number of stars they would give it; see Section 4.1) to its perceived social value, $V(s)$, is assumed to be an identity function. The presentational utility $U_{\textrm{pre}}(u;\phi)$ is grounded on the pragmatic listener $L_{1}$’s inference about a first-order pragmatic speaker $S_{1}$, who solely trades off information goal and social goal. Mathematically, the presentational utility can be formalized as $\log P_{L_{1}}(\phi|u)$. This quantity captures the extent to which a pragmatic listener $L_{1}$ infers a specific value trade-off $\phi$ under their internal model of a first-order pragmatic speaker $S_{1}$, where $P_{L_{1}}(s,\phi|u)\propto P_{S_{1}}(u|s,\phi)P(s)P(\phi)$. In other words, $\phi$ is a trade-off that the speaker $S_{2}$ wants to project towards a lower-order pragmatic listener $L_{1}$. The utterance distributions of the first-order pragmatic speaker $S_{1}$ is as follows:
@@
The informativeness and the expected social value of an utterance $u$ are both a function of how the literal listener $L_{0}$ interprets utterances $P_{L_{0}}(s|u)$, which is grounded out on the literal semantics [[u]](s)[\![u]\!](s) with a prior over the states $s$ likely to be communicated, i.e. $P_{L_{0}}(s|u)\propto[\![u]\!](s)\cdot P(s)$.
@@ 3 Language model evaluation suites
We design two model suites for evaluation that cover a range of characteristics that are thought to have implications for LLMs’ ability to capture human-like value trade-offs (see Appendix Table 1).
3.1 Closed-source model suite
In the closed-source setting, we aim to understand the behavioral tendencies of widely-used black-box models and how their reasoning-optimized variants might be adapting LLM behaviors in everyday contexts where value alignment is critical (cf. Zhou et al., 2025; Huang et al., 2025; Jiang et al., 2025). We evaluate three degrees of reasoning in Anthropic, Google, and OpenAI’s models: a) models that do not explicitly use any additional chain-of-thought reasoning (Claude-Sonnet-3.7, Gemini-Flash-2.0, and ChatGPT-4o), and b) the low and medium effort reasoning modes of their reasoning counterparts (Claude-Sonnet-3.7, Gemini-2.5-Flash, o4-mini). For the Gemini and o4 models, these effort levels can be specified directly by the parameters “low” and “medium”, but for Claude-Sonnet-3.7, which instead uses a specific token count, we map these values to 1k tokens and 8k tokens, respectively, following the values indicated in the Gemini API documentation.
3.2 Open-source model suite
In the open-source setting, we seek to understand which factors influence model behavior after preference fine-tuning by systematically evaluating the effects of base model family, preference dataset, and alignment algorithm on the resulting value trade-offs. We consider all unique configurations of two 7B parameter base models (Qwen2.5-Instruct (Yang et al., 2024) and Llama-3.1-Instruct (Grattafiori et al., 2024), two feedback datasets (UltraFeedback and Anthropic HH-RLHF  (Bai et al., 2022a), and two learning algorithms (DPO (Rafailov et al., 2024) and PPO). In particular, we aimed to have very different characteristics captured by each of the feedback datasets and base models to aid the interpretation of the final results. For example, since UltraFeedback is a synthetic feedback dataset designed to optimize for instruction following, truthfulness, and honesty attributes, we hypothesized that this dataset would induce a stronger weighting on informational utility in models than HH-RLHF, which more strongly indexes on the harmlessness attribute that might be more strongly associated with social and presentational utilities.
For each configuration (8 total), we initialize from an instruction-tuned model, perform one epoch of supervised fine-tuning (SFT) on the ‘chosen’ responses, and follow with one epoch of preference optimization using either DPO or PPO (implemented using OpenRLHF (Hu et al., 2024)) with ArmoRM (Wang et al., 2024) as the reward model. We evaluate each model’s behavior across evenly spaced checkpoints throughout the preference fine-tuning stage to trace the evolution of alignment and value trade-offs (see Section C.1 for full hyperparameter details).
@@ 4.1 Experimental vignettes
We provide models with the same set of vignettes given to human participants in Yoon et al. (2020), which describe socially sensitive situations in which a speaker must convey their judgement of a listener’s creation (e.g. a poem, presentation, cake, etc.). The speaker’s actual opinion, or true state $s$, is expressed on a scale from 1 to 5 stars, where 1 is the lowest or most negative opinion, and 5 is the highest. [^2] We deviate from the original paper’s 0-3 heart scale to provide LLMs with a scale that is most natural to their training data, particularly online reviews. We find that this 1-5 star scale captures the semantic range of the available utterance options better than the original 0-3 scale. We present models with the set of eight utterance options $u$ (four descriptor words and their negations) in a multiple choice format, randomizing the order options for every query to prevent positional biases:
@@ 4.2 Inferring cognitive model parameters
We implemented the inference model in Stan (Carpenter et al., 2017), a probabilistic programming language, and used its default Hamiltonian Monte Carlo (No-U-Turn sampler, Hoffman et al. (2014)) to perform approximate inference of model parameters (see Section C.2 for further details).
@@ 5.2 Closed-source model suite
Figure 3 shows the results of fitting the responses of Anthropic, Gemini, and OpenAI’s language models across three reasoning budgets (none, low, and medium) to the second-order speaker model. We establish that our inferred parameter values generalize to a held-out test split of the closed-source model data that includes all combinations of reasoning budget, vignette framing, communicative goal, and model family via a posterior predictive check, and find that the average MSE from our inferred parameters is significantly lower than that of randomly sampled parameter values from a reasonable prior ($z$-test: $\mu_{\textrm{inferred}}=0.03$, $\mu_{\textrm{random}}=0.06$; z=12.49z=-12.49; $p<0.001$; see Appendix Figure 6).
We begin with analyzing models’ default behavior in the absence of any explicit communicative goals (black lines). The parameter values of $\phi$ for the first-order speaker in $S_{2}$ measures the relative mixture of informativeness and social utility that a speaker $S_{2}$ wishes the other person to infer about them. We find that across model families, reasoning variants display higher $\phi$ values—a higher projected informational utility than social utility—than their non-reasoning counterparts. A linear mixed-effects model predicting the posterior mean $\phi$ from degrees of reasoning effect333model formula: phi $\sim$ reasoning_effort + (1|llm_family) + (1|framing) (reference level: no reasoning) with random intercepts of model family and vignette framing suggested a significant effect of both low and medium reasoning effort compared to the no-reasoning counterpart for default model behaviors ($t$-test: $\beta_{\text{low}}=0.228,t=6.338,p<.001$; $\beta_{\text{medium}}=0.211,t=5.846,p<.001$). The difference of the inferred $\phi$ among models of low and medium reasoning effort was not significant ($p=0.627$).
@@
Interpretable effects of manipulating communicative goals
@@
Figure 5: Open-source LLM results. Inferred values of informational, social, and presentational utilities $\boldsymbol{\omega}$, and projected mixture of informational and social utilities $\phi$, according to a cognitive model for LLMs’ training checkpoints across the RLHF process. Line variants indicate different combinations of base model and feedback dataset; rows = alignment method. Error bars indicate 95% high density region averaged across results from three framing manipulations. We find the largest shifts in values within the first quarter of training, with persistent effects of the choice of base model and pretraining data, compared to feedback dataset or alignment method.
@@ 5.3 Open-source model suite
Figure 5 shows the training dynamics of two base open-source LLMs, Qwen2.5-7B-Instruct (blue) and Llama-3.1-8B-Instruct (red), aligned to the UltraFeedback (dashed line) and Anthropic HH-RLHF (solid line) datasets, via DPO (top row) and PPO (bottom row). Across the different inferred parameters, we observe a number of consistent patterns within combinations of model and dataset. Across both PPO and DPO and the two feedback datasets, Qwen-instruct shows a higher $\omega_{\text{inf}}$, but lower weighting of $\omega_{\text{pre}}$ than Llama-instruct. The differences between the models’ weighting of social utility $\omega_{\text{soc}}$ are less pronounced, but still present, with Qwen-instruct generally converging to a lower weighting of social utility than Llama-instruct. The projected weighting towards informational utility in Qwen-instruct’s $\phi$, as well as its higher $\omega_{\text{inf}}$ compared to Llama-instruct aligns with prior work highlighting Qwen’s superior performance in mathematical and reasoning tasks compared to Llama (Gandhi et al., 2025; Zeng et al., 2025).
@@
For most of the inferred parameters, we do not observe significant qualitative differences in the training dynamic patterns resulting from PPO vs. DPO alignment methods. However, for the parameter $\phi$, PPO does appear to pull all four model and feedback dataset configurations to a similar mean value (approx. 0.7). In contrast, in the DPO case, Qwen-instruct appears to quickly converge to a greater weighting of informational utility, with $\phi$ almost equal to 1 in the case of alignment to both feedback datasets, which Llama-instruct shows more of a balance towards social utility (though it is still primarily information-leaning). The relatively minor distinction between PPO and DPO in our results may be partly due to training both methods for only a single epoch, and the fact that the Armo-RM reward model used in PPO, was trained on subsets of the same UltraFeedback and Anthropic HH-RLHF datasets, further reducing divergence between the two approaches.
Convergent evidence from deep learning literature
Our method present a novel approach to understanding the high-level behavior of LLMs, but notably, recovers patterns of behavior that have been found elsewhere in the deep learning literature. We highlight those connections here, as evidence of the promise of cognitive models for understanding more than just human behavior. In general, we see that the largest shifts in utility values across all four parameters happen within the first quarter of training, consistent with earlier findings on rapid adaptation during RL post-training in mathematical domains (Zhao et al., 2025). We also observe that the choice of base model and pretraining data may have an outsized impact on the resulting weighting of utilities: feedback dataset only shifts the trajectory established by the base model, but does not cause their behavioral profiles to converge over the course of training. In concurrent work, Christian et al. (2026) localize a similar finding to the reward models themselves, tracing biases in values back to the pretrained models upon which they are built. Prior work has also emphasized the significance of the base model and its pretraining data in disentangling the origins of cognitive biases (Itzhak et al., 2025), though our use of a shared supervised fine-tuning (SFT) stage on the same preference datasets across all models may attenuate these differences. Future work testing techniques for co-training for stable value alignment while preserving task performance would also be beneficial: initial proposals include multi-objective or constraint-based training, retaining some alignment data or synthetic proxies of the polite-speech scenarios, or using lightweight adapters for task-specific finetuning.